# An efficient wavelet-based physics-informed neural networks for singularly perturbed problems

### v1

@@ 3 Results and discussions
In this section, we present the results obtained using the proposed W-PINN method over a handful of SPPs and compare them with the performance of conventional PINNs. For a fair comparison, all network parameters are kept the same for both and are tabulated along with the examples. We utilize the Adam optimizer [45], which combines two stochastic gradient descent algorithms: gradient descent with momentum and root mean squared propagation (RMSProp). RMSprop, also known as adaptive gradient descent, adapts the learning rate with iterations. Adam optimizer is memory efficient and is well suited for problems with high-dimensional parameter spaces.
@@
For implementation purposes, the pytorch framework is used. Training is done on the NVIDIA RTX A6000, which has 48GB of GPU memory and 128GB of RAM. Conventional PINN requires roughly four minutes to complete $10,000$ iterations with a network having eight hidden layers, each containing $200$ neurons for $10,000$ collocation points. In contrast, W-PINN requires about one minute, which is significantly less. One of the reasons is that, as mentioned earlier, W-PINN does not require automatic differentiation for derivatives of the network’s output with respect to input parameters. Another reason for faster convergence is that the wavelet space is more suitable for addressing problems with sharp singularities.
As the performance of a neural network is highly sensitive to the initialization of its parameters, we performed each experiment five times and discussed the best of the five here. We begin with a singularly perturbed linear advection-diffusion ODE, then proceed to nonlinear and coupled nonlinear ODEs. Using one of the nonlinear ODEs, we investigate how the network’s hyperparameters affect its performance and also compare it with the approach proposed in [41]. Finally, we tested W-PINN over the FitzHugh-Nagumo model and predator-prey model.
@@
Such models, called Friedrich’s boundary layer models, are often used to show how difficult it is to model viscous flow boundary layers [32]. It is impossible for PINNs to capture the singularity that happens when $\epsilon\rightarrow 0$. In [32], improved PINNs are used to solve boundary layer problems with singular perturbation. In which the domain of the problem is conceptually divided into inner and outer regions to improve the application of PINNs. While the method is promising, its computational complexity and implementation challenges limit its practicality. W-PINNs offer significant advantages over this method.
Table 1: (Example 3.1) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 8/10
Number of hidden layers 6
Neurons per layer 100
Learning rate $10^{-4}$
Number of iterations $2\times 10^{4}$
Number of collocation points $10^{3}$
Table 2: (Example 3.1) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Mexican Gaussian
$2^{-1}$ 1.6e-04 5.8e-04 4.1e-04
$2^{-5}$ 3.3e-02 1.9e-04 2.7e-04
$2^{-7}$ 8.4e-01 9.2e-04 5.3e-04
Figure 4: Comparision of PINN with W-PINN for Example 3.1.
The parameters of the neural network for the Example 3.1 are described in Table 2. The results have been presented in Table 2 and Figure 4. Specifically, Table 2 provides the relative $L_{2}$-error and compares the W-PINN with the conventional PINN. This table shows that when $\epsilon=2^{-1}$, there is no singularity near the origin, and PINN approximates the solution well. However, when $\epsilon$ decreases, the exact solution of the problem has a strong singularity near the origin. This means that the layer at the origin becomes very sharp, and PINN is unable to accurately capture the behavior of the solution. W-PINN effectively resolves the singularity, with the approximate solution closely following the exact solution. Similar observations are drawn from Figure 4. It is shown in this figure that for $\epsilon=2^{-7}$, the W-PINN solution precisely matches the exact solution, whereas the PINN solution is far away from the exact one. Therefore, it can be concluded that for the problem, the W-PINN effectively handles the initial singularity effectively.
@@
Table 3: (Example 3.2) Network’s parameters.
@@
Table 4: (Example 3.2) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Sigmoid Gaussian Mexican Gaussian
activation activation
$2^{-4}$ 1.2e-04 2.1e-04 1.8e-04 2.3e-04
$2^{-7}$ 8.5e-02 2.2e-02 3.7e-04 1.3e-04
$2^{-10}$ 1.4 0.34 4.5e-03 9.6e-05
Figure 5: Comparision of PINN with W-PINN for Example 3.2.
@@
This example corresponds to a highly singularly perturbed non-linear equation with a known solution. Such an equation is a well-known tool that is frequently exploited to analyze problems in fluid dynamics. for Example 3.2, the parameters of the network are given in Table 4. The results are presented in Table 4 and Figure 5. We compare the W-PINN with the traditional PINN, and the PINN proposed in [41] (using Gaussian wavelet as an activation function) for various $\epsilon$ by considering the relative $L_{2}$-error in Table 4. As the singularity becomes more prominent, Table 4 illustrates that the W-PINN using Gaussian wavelet resolves the singularity to a greater extent.
The error obtained using PINN is very large for $\epsilon=2^{-10}$; in this case, PINNs are unable to detect the singularity, and also PINNs with the Gaussian wavelet as activation function [41] couldn’t resolve singularity satisfactorily, whereas W-PINN provides less error and effectively handles the singularity. One can get corresponding observations from Figure 5, in which solutions using PINN and W-PINN are plotted for various $\epsilon$.
@@
In addition, we investigate the relationship between the hyperparameters of W-PINN and its performance for this example. Figure 7 illustrates the relationship between the resolution level and the relative error. Here, we employ a network with $8$ hidden layers, each containing $200$ neurons, and fix $10,000$ collocation points. It shows that W-PINN gives optimal results for a specific resolution range (not too high nor too low). Lower resolution is incapable of capturing sharp singularities, and a large $J$ makes the loss function too complex to be optimized. The optimal range of the resolution is also dependent on the wavelet family. In this specific case, the W-PINN with Gaussian performed well for $J$, which is between $8$ and $14$, while this range is narrower for the Mexican hat. However, W-PINN performance also depends on the kind of wavelet used for a particular problem.
Figure 7 presents the effect of the number of collocation points on performance. We consider the identical network as previously used, with the Gaussian resolutions ($J_{G}=12$) and the Mexican hat resolutions ($J_{M}=11$). It is evident that the relative $L_{2}$-error decreases as the number of collocation points increases. However, the error decrease is not significant beyond a certain threshold. After a certain number of collocation points, additional points do not come with any new information about the behavior of the problem in consideration.
Figure 8 presents experiments on the trainable part of W-PINN. We employ $10,000$ collocation points for both Gaussian and Mexican, with resolutions of 12 and 11, respectively. It implies that a shallow network is incapable of learning the behaviors of the problem, regardless of the number of neurons per layer. However, a network with a significant number of layers and sufficient neurons per layer performs adequately. It is redundant to employ a large number of layers and neurons, which results in an increase in computational cost rather than accuracy.
@@
Figure 6: Resolution levels $(J)$ versus relative $L_{2}$-error.
@@
Figure 7: Number of collocation points $(N)$ versus relative $L_{2}$-error.
@@
Figure 8: Impact of various hidden layers and number of neurons per layer over relative $L_{2}$-error.
@@
Table 6 represents the parameters of the network for the example 3.3. The results have been displayed in Table 6 and Figure 9. Specifically, the table 6 compares the performance of the W-PINN and traditional PINN in terms of relative $L_{2}$-error across various $\epsilon$ values. By analyzing this table, it is evident that as the singularity becomes more prominent, our W-PINN effectively addresses the singularity to a greater degree compared to the PINN. Furthermore, Figure 9 is plotted for various $\epsilon$ values. This figure illustrates that as $\epsilon$ decreases, PINN encounters difficulties in capturing these singularities, as it is unable to incorporate them into its solutions. However, W-PINN efficiently adapts these singularities.
Table 5: (Example 3.3) Network’s parameters.
@@
Table 6: (Example 3.3) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Mexican Gaussian
$2^{-4}$ 8.6e-05 1.4e-04 4.7e-05
$2^{-7}$ 1.8e-03 8.3e-05 8.9e-06
$2^{-10}$ 5.3e-03 9.9e-05 6.5e-06
Figure 9: Comparision of PINN with W-PINN for Example 3.3.
Example 3.4.
Consider the following singularly perturbed non-linear coupled problem [49]
$\begin{cases}\epsilon_{1}\dfrac{du_{1}}{dx}+3u_{1}-\dfrac{1}{4}\exp(-u_{1}^{2})-u_{2}-x^{2}+1=0,~{}x\in(0,1],\\[6.0pt] \epsilon_{2}\dfrac{du_{2}}{dx}+4u_{2}-\cos(u_{2})-u_{1}-\exp(x)=0,~{}x\in(0,1],\\[6.0pt] u_{1}(0)=1,~{}u_{2}(0)=1.\end{cases}$ (8)
This example represents a system of non-linear singularly perturbed differential equations and has an unknown solution. In particular, we investigate three cases: The case-1 occurs when there is no singularity in the solution $u_{1}$, but the solution $u_{2}$ contains an abrupt singularity, i.e., $\epsilon_{1}=1$ and $\epsilon_{2}=2^{-10}$. The case-2 is the other way around; such a model allows us to study how fast and slow processes interact and influence the overall system behavior. The case-3 refers to singularity in solutions $u_{1}$ and $u_{2}$, i.e., $\epsilon_{1}=\epsilon_{2}=2^{-10}$. This case represents scenarios where both processes occur simultaneously and rapidly impact the overall behavior of the system.
In our experiment for the Example 3.4, we use the parameters for the network in Table 8. Table 8 and Figure 10 represent the solutions. Specifically, we present the relative $L_{2}$-error with the W-PINN and the traditional PINN for various $\epsilon_{1}$ and $\epsilon_{2}$ in Table 8. From this table, we also observe that the error is much less for W-PINN compared to PINN. The similar performance difference between PINN and W-PINN is evident in Figure 10.
Table 7: (Example 3.4) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 10/12
Number of hidden layers 8
Neurons per layer 200
Learning rate $10^{-5}$
Number of iterations $2\times 10^{4}$
Number of collocation points $10^{4}$
Table 8: (Example 3.4) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon_{1}$
$\epsilon_{2}$ Mexican Gaussian
111~{}~{}~{}~{}1 $u_{1}$ 9.8e-03 2.8e-04 1.1e-04
$2^{-10}$ $u_{2}$ 1.2e-02 5.2e-03 6.7e-04
$2^{-10}$ $u_{1}$ 7.1e-01 8.1e-03 5.7e-03
111~{}~{}~{}~{}1 $u_{2}$ 3.2e-02 5.1e-04 1.8e-04
$2^{-10}$ $u_{1}$ 5.4e-01 7.1e-03 6.8e-03
$2^{-10}$ $u_{2}$ 4.9e-02 1.1e-03 8.9e-04
Figure 10: Comparision of PINN with W-PINN for Example 3.4.
Table 9: (Example 3.5) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 11/12
Number of hidden layers 10
Neurons per layer 300
Learning rate $10^{-5}$
Number of iterations $4\times 10^{4}$
Number of collocation points $10^{4}$
Table 10: (Example 3.5) Relative $L_{2}$-error.
PINN W-PINN
$\tau~{}~{}$ Mexican Gaussian
$2^{-10}$ $v$ 5.9e-04 4.4e-03 9.2e-05
$w$   8.1e-02 8.2e-03 5.7e-04
Figure 11: Comparision of PINN with W-PINN for Example 3.5, using $\tau=2^{-10}$.
@@
For the numerical computations, we set: $a=1,~{}b=1,~{}I=0.1,~{}R=1,$ and $\tau=2^{-10}$, $v(0)=0.5$ and $w(0)=0.1$. The dynamics of the system are computed till $t=1$. Table 10 provides the parameters for this example 3.5. The results are presented in Table 10 and Figure 11. In particular, the relative $L_{2}$-error with the W-PINN and the traditional PINN is shown in Table 10. We present the W-PINN approximations of $v$ and $w$ in Figure 11. This figure is evident that an initial layer is present in the solution profile of $w$. Further, the figure illustrates that PINN fails to capture the sharp singularity of the solution and generates spurious oscillations near the singularity, whereas W-PINN effectively resolves the singularity.
Table 11: (Example 3.6) Network’s parameters.
@@
Level of resolution ($J_{M}/J_{G}$) 11/13
Number of hidden layers 8
@@
Table 12: (Example 3.6) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}~{}~{}~{}~{}~{}~{}$ Sigmoid Gaussian Mexican Gaussian
activation activation
$2^{-8}$  $x$ 0.342 5.8e-03 4.6e-03 6.3e-04
$y$ 0.997 0.991 1.8e-03 6.7e-04
Figure 12: Comparision of PINN with W-PINN for Example 3.6, using $\epsilon=2^{-8}$.

### v2

@@ 3 Results and discussions
In this section, we present the results obtained using the proposed method over a handful of examples and compare them with the performance of conventional PINNs and state-of-the-art methods. For a fair comparison, all network parameters are kept the same throughout and are tabulated along with the examples. We utilize the Adam optimizer [39], which combines two stochastic gradient descent algorithms: gradient descent with momentum and root mean squared propagation (RMSProp). RMSprop, also known as adaptive gradient descent, adapts the learning rate with iterations. Adam optimizer is memory efficient and is well suited for problems with high-dimensional parameter spaces.
@@
For implementation purposes, the PyTorch 2.4.1 with CUDA 12.1 is used. Training is done on the NVIDIA RTX A6000, which has 48GB of GPU memory . As the performance of a neural network is highly sensitive to the initialization of its parameters, so we perform each experiment five times and report relative $L_{2}$-error and average training time based on these five random runs.
Table 1: Parameters used for Example 3.1.
Parameters Value
Set of resolutions ($J_{M}/J_{G}$) [0,8]/[0,9]
Number of hidden layers 6
Neurons per layer 100
Number of collocation points $10^{3}$
Maximum number of iterations $2\times 10^{4}$
Table 2: Training time (in seconds) of PINN and W-PINN with network depth for Example 3.1.
Depth PINN W-PINN
2 58.3      21.4
4 80.1      26.2
8 131.3      33.4
16 223.4      41.7
32 367.1      58.8
@@
Such models, called Friedrich’s boundary layer models, are often used to show how difficult it is to model viscous flow boundary layers [22]. It is impossible for PINNs to capture the singularity that happens when $\epsilon\rightarrow 0$.
Table 3: Performance comparison of various methods for Example 3.1.
$\epsilon$ Method $L_{2}$-error
Average
Training Time
PINN 7.9e-05 1m 58s
Wavelet Activation [34] 3.9e-05 52s
$\epsilon=2^{-4}$ W-PINN (Gaussian) 2.5e-05 23s
W-PINN (Mexican hat) 8.1e-05 48s
PINN 0.83 -
Wavelet Activation [34] 0.85 -
$\epsilon=2^{-7}$ W-PINN (Gaussian) 5.3e-04 18s
W-PINN (Mexican hat) 9.2e-04 23s
PINN 0.84 -
Wavelet Activation [34] 0.85 -
$\epsilon=2^{-10}$ W-PINN (Gaussian) 3.1e-03 38s
W-PINN (Mexican hat) 4.2e-03 42s
Figure 2: Comparison of solutions obtained by PINN, PINN with wavelet activation and W-PINN methods for Example 3.1.
The parameters of the W-PINN for the Example 3.1 are listed in Table 2. Table 2 compares the training time of automatic differentiation based PINN and W-PINN. The experiments were conducted with $\epsilon=2^{-4}$ and each test was executed for $10^{3}$ iterations. The results demonstrate that conventional PINN exhibits a steeper growth in training time with increase in network depth, requiring up to $6.2$ times longer than W-PINN at depth $32$. This substantial difference in computational overhead can be attributed to W-PINN’s elimination of automatic differentiation calculations, which become increasingly expensive in conventional PINNs as network depth increases. The comparison of various methods have been presented in Table 3, which provides the relative $L_{2}$-error and average training time of the W-PINN, conventional PINN and PINN with gaussian-wavelet activation function [34]. This table shows the W-PINN approximates the solution with much smaller $L_{2}$-error for lower values of $\epsilon$. Similar observations are drawn from Figure 2, which demonstrate when $\epsilon=2^{-4}$, there is no such sharp singularity near the origin, all methods approximate the solution well. However, as $\epsilon$ decreases, the exact solution of the problem has a strong singularity near the origin. Then both PINN and PINN with wavelet activation function are unable to accurately capture the behavior of the solution, whereas W-PINN effectively resolves the singularity, with the predicted solution closely following the exact solution.
@@
Table 4: Parameters used for Example 3.2.
@@
Maximum number of iterations $2\times 10^{4}$
Table 5: Performance comparison of various methods for Example 3.2.
$\epsilon$ Method $L_{2}$-error
Average
Training Time
PINN 1.2e-04 31s
Wavelet Activation [34] 2.1e-04 58s
$\epsilon=2^{-4}$ W-PINN (Gaussian) 2.3e-04 48s
W-PINN (Mexican hat) 1.8e-04 51s
PINN 8.5e-02 1m 33s
Wavelet Activation [34] 2.2e-02 1m 14s
$\epsilon=2^{-7}$ W-PINN (Gaussian) 1.3e-04 21s
W-PINN (Mexican hat) 3.7e-04 24s
PINN 1.4 -
Wavelet Activation [34] 0.34 -
$\epsilon=2^{-10}$ W-PINN (Gaussian) 9.6e-05 27s
W-PINN (Mexican hat) 4.5e-04 48s
@@
This example corresponds to a highly singularly perturbed non-linear equation with a known solution. For Example 3.2, the parameters of the network are given in Table 4. We compare the performance of W-PINN with the traditional PINN, and the PINN proposed in [34] (using Gaussian wavelet as an activation function) for various $\epsilon$. The relative $L_{2}$-error and average training time for these are tabulated in Table 5, which illustrates that the W-PINN outperforms other methods as it takes much less training time and provides better approximation, especially for lower values of $\epsilon$. Figure 3 demonstrates similar observations, for $\epsilon=2^{-10}$ both conventional PINN and PINN with the Gaussian wavelet as activation function [34] are unable to detect the singularity satisfactorily, whereas W-PINN effectively handles the singularity.
@@
In addition, we investigate the relationship between the hyperparameters of W-PINN and its performance for this example. Figure 5 illustrates the relationship between the resolution level ($[0,J]$) and the relative $L_{2}$-error. Here, we employ a network with $8$ hidden layers, each containing $200$ neurons, and fix $10,000$ collocation points. It shows that W-PINN gives optimal results for a specific resolution range (not too high nor too low). Lower resolution is incapable of capturing sharp singularities, and a large $J$ makes the loss function too complex to be optimized. The optimal range of the resolution is also dependent on the wavelet family. In this specific case, the W-PINN with Gaussian performed well for $J$, which is between $8$ and $14$, while this range is narrower for the Mexican hat. However, W-PINN performance also depends on the kind of wavelet used for a particular problem.
Figure 3: Comparison of solutions obtained by PINN, PINN with wavelet activation and W-PINN methods for Example 3.2.
@@
Figure 4: Relative $L_{2}$-error variation with wavelet resolution levels $([0,J])$ for Gaussian and Mexican Hat wavelets.
@@
Figure 5: Relative $L_{2}$-error variation with number of collocation points $(N)$ for Gaussian and Mexican Hat wavelets.
@@
Figure 6: Impact of various hidden layers and number of neurons per layer on performance of W-PINN using Gaussian wavelet(left) and Mexican Hat wavelet(right).
Figure 5 presents the effect of the number of collocation points on performance. We consider the identical network as previously used, with the Gaussian resolutions ($J_{G}=[0,12]$) and the Mexican hat resolutions ($J_{M}=[0,11]$). It is evident that the relative $L_{2}$-error decreases as the number of collocation points increases. However, the error decrease is not significant beyond a certain threshold. After a certain number of collocation points, additional points do not come with any new information about the behavior of the problem in consideration.
Figure 6 presents experiments on the trainable part of W-PINN. We employ $10,000$ collocation points for both Gaussian and Mexican, with resolutions of $[0,12]$ and $[0,11]$, respectively. It implies that a shallow network is incapable of learning the behaviors of the problem, regardless of the number of neurons per layer. However, a network with a significant number of layers and sufficient neurons per layer performs adequately. It is redundant to employ a large number of layers and neurons, which results in an increase in computational cost rather than accuracy.
@@
Table 6: Parameters used for Example 3.3.
@@
Maximum number of iterations $10^{4}$
Table 7: Performance comparison of various methods for Example 3.3
$\epsilon$ Method $L_{2}$-error
Average
Training Time
PINN 8.6e-05 2m 26s
Wavelet Activation [34] 3.2e-04 4m 49s
$\epsilon=2^{-4}$ W-PINN (Gaussian) 4.7e-05 28s
W-PINN (Mexican hat) 1.4e-04 34s
PINN 1.8e-03 4m 41s
Wavelet Activation [34] 4.1e-03 6m 13s
$\epsilon=2^{-7}$ W-PINN (Gaussian) 8.9e-06 14s
W-PINN (Mexican hat) 8.3e-05 21s
PINN 5.3e-03 7m 14s
Wavelet Activation [34] 8.4e-03 12m 28s
$\epsilon=2^{-10}$ W-PINN (Gaussian) 6.5e-06 13s
W-PINN (Mexican hat) 9.9e-05 28s
Figure 7: Comparison of solutions obtained by PINN, PINN with wavelet activation and W-PINN methods for Example 3.3.
Table 6 represents the parameters of the network for the Example 3.3. The results have been displayed in Table 7 and Figure 7. Specifically, the Table 7 compares the performance of W-PINN, traditional PINN, and the PINN with gaussian wavelet as activation function [34] for various $\epsilon$ in terms of relative $L_{2}$-error and average training time. From this table, it is evident that W-PINN takes much less training time and approximates the solution to a better extent. Figure 7 demonstrates as the singularity becomes more prominent, W-PINN effectively addresses the singularity to a greater degree at both ends, in contrast, both PINNs method couldn’t resolve singularity and introduce unnatural oscillations over the domain.
@@
Table 8: Parameters used for Example 3.4.
@@
Set of resolutions ($J_{M}/J_{G}$) [0,10]/[0,11]
Number of hidden layers 10
@@
Number of iterations $2\times 10^{4}$
Table 9: Comparison of various methods for Example 3.4 with $\tau=0.15$.
Methods $L_{2}$-error       Average
     $(v/w)$ Training Time
PINN $045.9e-$ 04       6m 38s
$028.1e-$ 02
Wavelet Activation $032.1e-$ 03       8m 22s
$0.11$
W-PINN (Gaussian) $059.2e-$ 05       1m 06s
$045.7e-$ 04
W-PINN (Mexican hat) $034.4e-$ 03       1m 36s
$038.2e-$ 03
Figure 8: Comparison of solution of FHN model 3.4 obtained by PINN, PINN with wavelet activation and W-PINN methods using $\tau=2^{-10}$.
For the numerical computations, we set: $a=1,~{}b=1,~{}I=0.1,~{}R=1,$ and $\tau=2^{-10}$, $v(0)=0.5$ and $w(0)=0.1$. The dynamics of the system are computed till $t=1$. Table 8 provides the parameters for this Example 3.4. From Table 9, it is evident that W-PINN gets a more accurate prediction and takes significantly less training time. Figure 8 compares the W-PINN approximations of $v$ and $w$ with PINN and PINN with wavelet activation. The figure illustrates that a strong initial layer is present in the solution profile of $w$, that both PINN fail to capture and generate spurious oscillations near the singularity, whereas W-PINN effectively resolves the singularity.